OpenStudy (anonymous):

can anyone help me with this question: find the area for each circle with a radius of 2.2 cm I will award a medal

OpenStudy (anonymous):

This is very similar to the previous problem Area = pi(r)^2 r = 2.2 pi(2.2)^2 Can you solve from here?

OpenStudy (anonymous):

not really sure

OpenStudy (anonymous):

Well first you should square 2.2 2.2^2 = 4.84 This leaves you with 4.84pi Do you remember what I said the value of pi was?

OpenStudy (anonymous):

pi is 3.12

OpenStudy (anonymous):

so my answer would be 15.2

OpenStudy (anonymous):

3.14 actually :) You are going to need to remember that for the rest of your life. It will come up so often Multiply 4.84 with 3.14 (4.84)(3.14) = ?

OpenStudy (anonymous):

Correct Punkchick. The other guy was correct in the process except that he used 2.2^2 = 4.48 not 4.84